# Testing the limits of ITkPixV2: the ATLAS inner tracker pixel detector readout chip

### v1

@@ Abstract
The ITkPixV2 chip is the final production readout chip for the ATLAS Phase 2 Inner Tracker (ITk) upgrade at the upcoming High-Luminosity LHC (HL-LHC). Due to the extraordinarily high peak luminosity at the HL-LHC of $5\times 10^{34}$ cm-1s-1, ITkPixV2 must meet significant increases in nearly all design requirements, including a 10x increase in trigger rate, a 7.5x increase in hit rate, a 3x increase in radiation tolerance, and a 12.5x decrease in pixel current draw per unit area, all while maintaining a similar power per unit area as present pixel detectors. Here we present the first measurements of the ITkPixV2 chip operated at the limits of the full chip design requirements, including in particular a measurement of the activity-induced current of the chip as a function of increasing hit rate.
@@ 3 Experimental setup
[…]
@@
Figure 3: Hit/bit rates as a function of X-ray tube current, with linear fits to low-current regions extrapolated.
@@ 4 Results
Figure 4: Digital current as a function of hit rate.

### v3

@@ Abstract
The ITkPixV2 chip is the final production readout chip for the ATLAS Phase 2 Inner Tracker (ITk) upgrade at the upcoming High-Luminosity LHC (HL-LHC). Due to the extraordinarily high peak luminosity at the HL-LHC of $5\times 10^{34}$ cm-1s-1, ITkPixV2 must meet significant increases in nearly all design requirements compared to the current ATLAS Inner Detector (ID), including a 10x increase in trigger rate, a 7.5x increase in hit rate, a 3x increase in radiation tolerance, and a 12.5x decrease in pixel current draw per unit area, all while maintaining a similar power per unit area as present pixel detectors. Here we present the first measurements of the ITkPixV2 chip operated at the limits of the full chip design requirements, including in particular a measurement of the activity-induced current of the chip as a function of increasing hit rate.
@@ 3 Experimental setup
Chip bias voltage was set and monitored with a Keithley 2400 to an accuracy of 0.012% [10]. The chip itself carried a 100 $\mu$m thick sensor produced and bump bonded to the ASIC by Micron Technology Inc., with a bias voltage of 35V applied in all tests. The chip was mounted on a cooling chuck operated at 10 C, and the on-chip NTC temperature reading was approximately 18 C throughout.
@@
Figure 3: Hit/bit rates as a function of X-ray tube current. Linear fits are shown, with slope $m$ defined fitted slope in the region with hit rate $<$ 1 GHz/cm2. These fits are extrapolated to show the relative hit loss at high X-ray tube current values.
@@ 4 Results
Figure 4: Digital current as a function of hit rate. Linear fits are shown with slope $m$ corresponding to the fitted slope in the region before the “knee:" 2 GHz/cm2 for the 4-lane configuration, and 0.45 GHz/cm2 for the 1-lane configuration.